# Assessing Small Language Models for Code Generation: An Empirical Study with Benchmarks

### v3

@@ 3.1 SLMs Performance (RQ1)
Besides functional correctness, it is important to understand how consistently each model performs, which we examine through a stability score introduced in section 2.6. Table 4 presents the normalised stability scores of all the selected SLMs across different pass@k levels, alongside their standard deviation and coefficient of variation. These metrics reflect the stability of each model’s performance across the three benchmarks. Figure 3 offer a visual breakdown of these metrics. It shows the standard deviation and coefficient of variation for each model at different pass@k levels. In the figure, the models are plotted along the x-axis, while the values of standard deviation and coefficient of variation are shown on the y-axis. This visualisation highlights how much each model’s performance fluctuates from its mean across benchmarks. Models with lower standard deviation and coefficient of variation, along with a smaller gap between the two, are more stable in their performance across benchmarks. Figure 4 showcases the rank@k position of the models based on their stability scores across all pass@k levels. These results across the table and figures highlight the following insights about SLMs stability. The ranking indicates models from Group 3 generally achieve high normalised stability scores. For example, Qwen2.5-Coder 7B consistently ranks first across all pass@k levels with a Stability Score of 1.00, while other Group 3 models such as OpenCodeInterpreter 6.7B, Yi-Coder 9.0B, and CodeGemma 7B also obtain high scores, typically above 0.30. In Group 2, Stability Scores vary widely. Qwen2.5-Coder 3.0B achieves a Stability Score of 0.40 and ranks among the top five models, whereas PolyCoder 2.7B and CodeGemma 2.0B record low scores, close to 0.07 and 0.01, respectively. Models in Group 1 also show mixed results. Despite the small size, OpenCodeInterpreter 1.3B and Qwen2.5-Coder 1.5B achieve moderate Stability Scores (e.g., 0.43 and 0.25, respectively) and are ranked within the top 10 across multiple pass@k levels. In contrast, PolyCoder 0.4B and Incoder 1.3B obtain some of the lowest Stability Scores, with values as low as 0.00 and 0.17.
@@
The observed performance of the models was further validated statistically by conducting a two-way ANOVA test. This test assesses any influence on performance by (i) the size of the model (group) and (ii) the benchmark used, as well as whether there is any interaction between these factors. As reported in Table 5, the results show that the model group has a statistically significant effect on performance across all pass@k levels (p 0.05). In contrast, the choice of benchmark and the interaction between model group and benchmark were not statistically significant. To further understand where the differences lie, Table 6 reports the Tukey’s HSD post-hoc test results. The results reveal that Group 3 models (largest) significantly outperform Group 2 models (mid-sized) at pass@1, while the differences at pass@2, pass@5, and pass@10 are marginally significant or lie near the statistical threshold. However, the performance gap between Group 1 (smallest) and Group 3 is not statistically significant in pass@1,2, and 5 but significant in pass@10. Similarly, the performance gap between groups 1 and 2 is not statistically significant. However, their mean difference is negative in all pass@k.
@@ 3.2 Performance Trade-Offs (RQ2)
| # | Model Name | Mean_pass@1 | Mean_pass@2 | Mean_pass@5 | Mean_pass@10 | Average VRAM Usage (GB) | Average Inference Time (Sec) |
| Group 1: SLMs parameter size 1.5B | |||||||
| 1 | PolyCoder_0.4B | 0.01 | 0.02 | 0.02 | 0.02 | 5.03 | 2.49 |
| 2 | Incoder_1.3B | 0.06 | 0.08 | 0.11 | 0.13 | 13.88 | 3.06 |
| 3 | Phi1_1.3B | 0.41 | 0.45 | 0.49 | 0.50 | 8.45 | 1.28 |
| 4 | DeepSeek-Coder_1.3B | 0.43 | 0.48 | 0.54 | 0.57 | 12.12 | 1.19 |
| 5 | OpenCodeInterpreter_1.3B | 0.53 | 0.58 | 0.62 | 0.64 | 11.97 | 1.06 |
| 6 | Yi-Coder_1.5B | 0.49 | 0.54 | 0.58 | 0.61 | 11.99 | 0.80 |
| 7 | Qwen2.5-Coder_1.5B | 0.54 | 0.60 | 0.65 | 0.68 | 6.55 | 1.39 |
| Group 2: SLMs parameter size 1.5B to 3B | |||||||
| 8 | CodeGemma_2.0B | 0.28 | 0.34 | 0.39 | 0.42 | 8.70 | 1.78 |
| 9 | PolyCoder_2.7B | 0.03 | 0.04 | 0.05 | 0.07 | 14.15 | 5.58 |
| 10 | Stable Code_3.0B | 0.27 | 0.34 | 0.42 | 0.46 | 17.57 | 1.64 |
| 11 | StarCoder2_3.0B | 0.42 | 0.48 | 0.53 | 0.57 | 7.99 | 1.03 |
| 12 | Qwen2.5-Coder_3.0B | 0.59 | 0.64 | 0.69 | 0.72 | 10.77 | 1.69 |
| Group 3: SLMs parameter size 3B to 10B | |||||||
| 13 | Incoder_6.7B | 0.10 | 0.13 | 0.17 | 0.19 | 23.53 | 7.74 |
| 14 | DeepSeek-Coder_6.7B | 0.56 | 0.62 | 0.67 | 0.70 | 23.30 | 3.34 |
| 15 | OpenCodeInterpreter_6.7B | 0.67 | 0.70 | 0.73 | 0.74 | 23.32 | 2.67 |
| 16 | CodeLlama_7.0B | 0.37 | 0.44 | 0.51 | 0.55 | 23.58 | 2.11 |
| 17 | StarCoder2_7.0B | 0.47 | 0.52 | 0.57 | 0.60 | 17.00 | 2.92 |
| 18 | CodeGemma_7.0B | 0.49 | 0.59 | 0.68 | 0.74 | 23.38 | 1.67 |
| 19 | Qwen2.5-Coder_7.0B | 0.65 | 0.70 | 0.74 | 0.77 | 23.70 | 0.97 |
| 20 | Yi-Coder_9.0B | 0.59 | 0.65 | 0.71 | 0.73 | 20.33 | 3.99 |
@@ 3.3 Multilingual Consistency (RQ3)
We examine the multilingual performance of the models on two multilingual benchmarks: HumanEvalPack and CodeXGLUE. These benchmarks evaluate functional correctness and lexical similarity across multiple programming languages, providing insight into how model performance varies across different languages. Figure 6 presents the distribution of pass@k scores across four languages—Python, C++, Java, and JavaScript—in HumanEvalPack. The box plots display the median, interquartile range, and outliers for each language. At pass@1, Java achieves the highest median score (0.25), followed by Python and JavaScript (both 0.23), and C++ (0.16). This relative ranking remains consistent at higher k-values. For instance, at pass@10, Java reaches a median of 0.40, compared to 0.33 for Python, 0.35 for JavaScript, and 0.26 for C++. C++ consistently exhibits the lowest performance, particularly with narrow interquartile ranges and low upper bounds, indicating limited variation and poor model accuracy for this language. As there are observable differences in model performance across programming languages, we conducted a one-way ANOVA test to evaluate whether these differences are statistically significant. Table 10 presents the results of a one-way ANOVA test conducted on the pass@k scores obtained from the HumanEvalPack benchmark. The p-values are 0.46 for pass@1, 0.39 for pass@2, 0.32 for pass@5, and 0.27 for pass@10. These values are greater than 0.05, indicating that there is no statistically significant difference in model performance across the four programming languages evaluated.
In addition to pass@k based functional correctness, we also analyse natural language generation quality using BLEU scores on the CodeXGLUE benchmark. Figure 7 shows how frequently each language appears among the top in BLEU score ranks across the 20 models. Python appears in Rank 1 for 14 models and Rank 2 for 6 models, making it the most dominant language in high BLEU score positions. PHP ranks first in 6 models and second in 13, indicating strong performance among models in lower-ranked positions as well. Java dominates Rank 3, appearing 15 times, while Go appears most frequently in Rank 4 with 12 instances. Ruby appears 11 times in Rank 6 and 6 times in Rank 5, showing a strong presence in lower ranks. In contrast, JavaScript appears only 3 times in Rank 4, 10 times in Rank 5 and 7 times in Rank 6. It does not feature at all in the top two positions. These results highlight a clear skew toward Python and PHP in BLEU score rankings, while Java, Go, Ruby, and JavaScript frequently occupy mid-to-lower ranks.
@@ 4.1 Model Size on Code Generation Quality (RQ1)
This observation aligns with prior empirical studies. For example, Codex by Chen et al. [6] shows a strong correlation between model size and pass@k accuracy in code generation. Peng et al. [19] similarly observed minimal variation in benchmark performance compared to variation due to model size, suggesting that parameter scaling is a dominant driver of effectiveness. In addition, Ye et al. [20] emphasised that generalisation across code tasks becomes robust once a model crosses a certain size threshold, with benchmark differences becoming negligible. These prior findings highlight a reliable cross-task behaviour of SLMs, especially when evaluated at scale. The lack of benchmark sensitivity suggests that improvements in model design and training translate well across tasks with differing structural and semantic requirements. It further implies that benchmark-agnostic development and evaluation strategies may be valid for SLMs, particularly when considering architectural scaling or transfer learning.
@@
Performance and Stability of Large vs. Smaller SLMs: The obtained results of our evaluation show that larger SLMs tend to outperform smaller SLMs in both functional correctness and stability. They generally occupy top positions across evaluation metrics, indicating strong performance and consistent behaviour across diverse code generation tasks. However, our data also reveal that this trend does not always hold true; for example, some small and mid-sized models (e.g., OpenCodeInterpreter 1.3B, Qwen2.5-Coder 3.0) challenge it. Some models from lower parameter groups consistently secure high rankings and demonstrate notable performance and stability. These findings suggest that while scaling up the number of parameters generally improves models performance, it is not the only way to achieve strong performance in code generation. Well-designed and trained smaller SLMs can achieve competitive results when optimised effectively.
@@
Similar patterns have been documented in earlier empirical work. For example, Ye et al. [20] observed that once SLMs surpass a certain parameter threshold, additional gains in functional correctness become marginal. This resonates with our finding that while larger models generally perform well, several smaller and mid-sized models also achieve high accuracy and stability, suggesting diminishing returns from parameter scaling alone. It indicates that at some point, it becomes more beneficial to focus on how the model is trained or what data it utilises, rather than merely increasing its size. Similarly, Xu et al. [1] and Wang et al. [2] emphasised the feasibility of achieving competitive performance from well-optimised SLMs, especially in resource-constrained environments. Our results support this by demonstrating that small models, such as Qwen2.5-Coder 3.0, can outperform several larger models. Furthermore, Subramanian et al. [3] emphasised the underappreciated potential of SLMs in real-world Software Engineering workflows, calling for more empirical performance assessments that go beyond size alone. Our evaluation aligns with this perspective by showing that model size alone is not a reliable predictor of practical utility or performance consistency, especially in diverse code generation scenarios.

### v4

@@ 3.1 SLMs Performance (RQ1)
Besides functional correctness, it is important to understand how consistently each model performs, which we examine through a stability score introduced in section 2.6. Table 6 presents the normalised stability scores of all the selected SLMs across different pass@k levels, alongside their standard deviation and coefficient of variation. These metrics reflect the stability of each model’s performance across the three benchmarks. Figure 2 offer a visual breakdown of these metrics. It shows the standard deviation and coefficient of variation for each model at different pass@k levels. In the figure, the models are plotted along the x-axis, while the values of standard deviation and coefficient of variation are shown on the y-axis. This visualisation highlights how much each model’s performance fluctuates from its mean across benchmarks in different pass@k levels. Models with lower standard deviation and coefficient of variation, along with a smaller gap between the two, are more stable in their performance across benchmarks. Table 7 showcases the rank@k position of the models based on their normalised stability scores across all pass@k levels. These results across the table and figures highlight the following insights about SLMs stability.
Overall, Group 3 models show the highest stability. For example, Qwen2.5-Coder 7B achieves the maximum Stability Score of 1.00 across all pass@k levels. Other Group 3 models, such as OpenCodeInterpreter 6.7B, Yi-Coder 9.0B, also obtain decent Stability Scores of 0.52 and 0.37, respectively, at the pass@1 level. In contrast, Groups 1 and 2 exhibit wider variation. While Qwen2.5-Coder 3.0B (0.40 at pass@1) and OpenCodeInterpreter 1.3B (0.43 at pass@1) rank within the top ten across multiple pass@k levels, models like PolyCoder 0.4B and Incoder 1.3B obtain some of the lowest Stability Scores, with values as low as 0.00 and 0.17.
However, the stability results also reveal several nuanced patterns that are not captured by functional correctness alone. Many Group 3 models become more stable as "k(e.g., 1,2,5,10)" increases; for instance, CodeGemma 7B and StarCoder2 7B show noticeably lower variability at pass@10 than at pass@1, suggesting that larger models benefit from multi-sample generation where randomness is averaged out. At the same time, stability does not always correlate with performance. Some models with competitive pass@k scores exhibit comparatively high fluctuation—StarCoder2 7B, for example, achieves a respectable mean pass@1 of 0.47 but has a modest Stability Score of 0.19, indicating inconsistent behaviour across benchmarks. Conversely, certain smaller models demonstrate both strong performance and high stability. OpenCodeInterpreter 1.3B, despite its size, attains a mean pass@1 of 0.53 alongside an NSS of 0.43, outperforming several larger models. A notable pattern emerges within model families as well: the Qwen family demonstrates consistently high stability relative to size, particularly Qwen2.5-Coder 3B and Qwen2.5-Coder 7B, which maintain low variation and strong NSS values across all pass@k levels. This suggests that architectural and training pipeline design can play a substantial role in promoting stability, independent of parameter count.
The observed performance of the models was further validated statistically by conducting a two-way ANOVA test. This test assesses any influence on performance by (i) the size of the model (group) and (ii) the benchmark used, as well as whether there is any interaction between these factors. As reported in Table 8, the results show that the model group has a statistically significant effect on performance across all pass@k levels (p 0.05). In contrast, the choice of benchmark and the interaction between model group and benchmark were not statistically significant.
To further understand where the differences lie, Table 9 reports the Tukey’s HSD post-hoc test results. The results reveal that Group 3 models (largest) significantly outperform Group 2 models (mid-sized) at pass@1, while the differences at pass@2, pass@5, and pass@10 are marginally significant or lie near the statistical threshold. However, the performance gap between Group 1 (smallest) and Group 3 is not statistically significant in pass@1,2, and 5 but significant in pass@10. Similarly, the performance gap between groups 1 and 2 is not statistically significant. However, their mean difference is negative in all pass@k.
@@ 3.2 Performance Trade-Offs (RQ2)
| Model_Parameter | Mean_pass@1 | Mean_pass@2 | Mean_pass@5 | Mean_pass@10 | Avg VRAM (GB) | Avg Time (Sec) |
| Group 1: SLMs parameter size 1.5B | ||||||
| PolyCoder_0.4B | 0.01 | 0.02 | 0.02 | 0.02 | 5.03 | 2.49 |
| Incoder_1.3B | 0.06 | 0.08 | 0.11 | 0.13 | 13.88 | 3.06 |
| Phi1_1.3B | 0.41 | 0.45 | 0.49 | 0.50 | 8.45 | 1.28 |
| DeepSeek-Coder_1.3B | 0.43 | 0.48 | 0.54 | 0.57 | 12.12 | 1.19 |
| OpenCodeInterpreter_1.3B | 0.53 | 0.58 | 0.62 | 0.64 | 11.97 | 1.06 |
| Yi-Coder_1.5B | 0.49 | 0.54 | 0.58 | 0.61 | 11.99 | 0.80 |
| Qwen2.5-Coder_1.5B | 0.54 | 0.60 | 0.65 | 0.68 | 6.55 | 1.39 |
| Group 2: SLMs parameter size 1.5B to 3B | ||||||
| CodeGemma_2.0B | 0.28 | 0.34 | 0.39 | 0.42 | 8.70 | 1.78 |
| PolyCoder_2.7B | 0.03 | 0.04 | 0.05 | 0.07 | 14.15 | 5.58 |
| Stable Code_3.0B | 0.27 | 0.34 | 0.42 | 0.46 | 17.57 | 1.64 |
| StarCoder2_3.0B | 0.42 | 0.48 | 0.53 | 0.57 | 7.99 | 1.03 |
| Qwen2.5-Coder_3.0B | 0.59 | 0.64 | 0.69 | 0.72 | 10.77 | 1.69 |
| Group 3: SLMs parameter size 3B to 10B | ||||||
| Incoder_6.7B | 0.10 | 0.13 | 0.17 | 0.19 | 23.53 | 7.74 |
| DeepSeek-Coder_6.7B | 0.56 | 0.62 | 0.67 | 0.70 | 23.30 | 3.34 |
| OpenCodeInterpreter_6.7B | 0.67 | 0.70 | 0.73 | 0.74 | 23.32 | 2.67 |
| CodeLlama_7.0B | 0.37 | 0.44 | 0.51 | 0.55 | 23.58 | 2.11 |
| StarCoder2_7.0B | 0.47 | 0.52 | 0.57 | 0.60 | 17.00 | 2.92 |
| CodeGemma_7.0B | 0.49 | 0.59 | 0.68 | 0.74 | 23.38 | 1.67 |
| Qwen2.5-Coder_7.0B | 0.65 | 0.70 | 0.74 | 0.77 | 23.70 | 0.97 |
| Yi-Coder_9.0B | 0.59 | 0.65 | 0.71 | 0.73 | 20.33 | 3.99 |
Beyond these broader trends, several additional patterns emerge when examining the results more closely. Some models exhibit poor efficiency relative to their size, for example, PolyCoder 2.7B requires 14.15 GB of VRAM and has the slowest inference time in the entire study (5.58 seconds), yet achieves only 0.03 pass@1. Similarly, CodeLlama 7B consumes 23.58 GB of VRAM, comparable to other Group 3 models, but delivers only 0.37 pass@1, making it one of the least efficient models in terms of accuracy per memory cost. Conversely, several small and mid-sized models demonstrate highly favorable trade-offs. Phi-1 1.3B achieves 0.41 pass@1 with only 8.45 GB VRAM and a 1.28 second inference time, outperforming multiple larger models, while StarCoder2 3B delivers 0.42 pass@1 with just 7.99 GB VRAM and a 1.03 second inference time. These comparisons highlight that parameter size alone does not predict efficiency and that well-optimized architectures and training pipelines can allow smaller SLMs to rival or surpass larger ones in practical performance.
@@ 3.3 Multilingual Consistency (RQ3)
We examine the multilingual performance of the models on two multilingual benchmarks: HumanEvalPack and CodeXGLUE. These benchmarks evaluate functional correctness and lexical similarity across multiple programming languages, providing insight into how model performance varies across different languages. Figure 4 presents the distribution of pass@k scores across four languages, i.e., Python, C++, Java, and JavaScript, in HumanEvalPack. The box plots display the median, interquartile range, and outliers for each language. At pass@1, Java achieves the highest median score (0.25), followed by Python and JavaScript (both 0.23), and C++ (0.16). This relative ranking remains consistent at higher k-values. For instance, at pass@10, Java reaches a median of 0.40, compared to 0.33 for Python, 0.35 for JavaScript, and 0.26 for C++. C++ consistently shows the lowest performance, particularly with narrow interquartile ranges and low upper bounds, indicating limited variation and poor model accuracy for this language. This behaviour aligns with the fact that C++ problems often involve stricter type constraints and more complex syntax than other languages, which increases the difficulty of generating functionally correct code. As there are observable differences in model performance across programming languages, we conducted a one-way ANOVA test to evaluate whether these differences are statistically significant. Table 13 presents the results of a one-way ANOVA test conducted on the pass@k scores obtained from the HumanEvalPack benchmark. The p-values are 0.46 for pass@1, 0.39 for pass@2, 0.32 for pass@5, and 0.27 for pass@10. These values are greater than 0.05, indicating that there is no statistically significant difference in model performance across the four programming languages evaluated.
In addition to pass@k based functional correctness, we also analyse natural language generation quality using BLEU scores on the CodeXGLUE benchmark. Figure 5 shows how frequently each language appears among the top in BLEU score ranks across the 20 models. Python appears in Rank 1 for 14 models and Rank 2 for 6 models, making it the most dominant language in high BLEU score positions. PHP ranks first in 6 models and second in 13, indicating strong performance among models in lower-ranked positions as well. Java dominates Rank 3, appearing 15 times, while Go appears most frequently in Rank 4 with 12 instances. Ruby appears 11 times in Rank 6 and 6 times in Rank 5, showing a strong presence in lower ranks. In contrast, JavaScript appears only 3 times in Rank 4, 10 times in Rank 5 and 7 times in Rank 6. It does not feature at all in the top two positions. These results highlight a clear skew toward Python and PHP in BLEU score rankings, while Java, Go, Ruby, and JavaScript frequently occupy mid-to-lower ranks. Python and PHP achieve higher BLEU ranks because their documentation styles in CodeXGLUE are concise, making them easier for SLMs to match lexically. In contrast, languages like Java, Go, Ruby, and JavaScript use more variable documentation patterns, which naturally lowers lexical overlap. Since BLEU measures surface similarity rather than semantic quality, these differences reflect documentation style and training-data distribution rather than inherent model weakness in those languages.
@@ 4.1 Model Size on Code Generation Quality (RQ1)
This observation aligns with prior empirical studies. For example, Codex by Chen et al.[Chen2021EvaluatingLL] shows a strong correlation between model size and pass@k accuracy in code generation. Peng et al.[peng2024humaneval] similarly observed minimal variation in benchmark performance compared to variation due to model size, suggesting that parameter scaling is a dominant driver of effectiveness. In addition, Ye et al.[ye2024benchmarking] emphasised that generalisation across code tasks becomes robust once a model crosses a certain size threshold, with benchmark differences becoming negligible. These prior findings highlight a reliable cross-task behaviour of SLMs, especially when evaluated at scale. The lack of benchmark sensitivity suggests that improvements in model design and training translate well across tasks with differing structural and semantic requirements. It further implies that benchmark-agnostic development and evaluation strategies may be valid for SLMs, particularly when considering architectural scaling or transfer learning.
@@
Performance and Stability of Large vs. Smaller SLMs: The obtained results of our evaluation show that larger SLMs tend to outperform smaller SLMs in both functional correctness and stability. They generally occupy top positions across evaluation metrics, indicating strong performance and consistent behaviour across diverse code generation tasks. However, our data also reveal that this trend does not always hold true; for example, some small and mid-sized models (e.g., OpenCodeInterpreter 1.3B, Qwen2.5-Coder 3B) challenge it. Some models from lower parameter groups consistently secure high rankings and demonstrate notable performance and stability. These findings suggest that while scaling up the number of parameters generally improves models’ performance, it is not the only way to achieve strong performance in code generation. Well-designed and trained smaller SLMs can achieve competitive results when optimised effectively.
@@
Similar patterns have been documented in earlier empirical work. For example, Ye et al.[ye2024benchmarking] observed that once SLMs surpass a certain parameter threshold, additional gains in functional correctness become marginal. This resonates with our finding that while larger models generally perform well, several smaller and mid-sized models also achieve high accuracy and stability, suggesting diminishing returns from parameter scaling alone. It indicates that at some point, it becomes more beneficial to focus on how the model is trained or what data it utilises, rather than merely increasing its size. Similarly, Xu et al.[xu2024survey] and Wang et al.[Wang2024ACS] emphasised the feasibility of achieving competitive performance from well-optimised SLMs, especially in resource-constrained environments. Our results support this by demonstrating that small models, such as Qwen2.5-Coder 3.0, can outperform several larger models. Furthermore, Subramanian et al.[Subramanian2025survey] emphasised the underappreciated potential of SLMs in real-world Software Engineering workflows, calling for more empirical performance assessments that go beyond size alone. Our evaluation aligns with this perspective by showing that model size alone is not a reliable predictor of practical utility or performance consistency, especially in diverse code generation scenarios.
Although this study focuses on SLMs in the 0.4B–10B parameter range, it is useful to briefly situate these findings with LLMs. Previous studies by Zheng et al. [zheng2024well] report results for GPT-3.5 on the HumanEval benchmark using settings such as temperature = 0.4 and a single sample (n = 1) per problem. On the other hand, Li and Murr [xia2024top]evaluate recent GPT models on HumanEval in a zero-shot setup for n=10 samples via the OpenAI API, and Chen et al. [Chen2021EvaluatingLL] report HumanEval results for Codex with sample generation at temperature = 0.8 and up to n=100 samples per problem. However, these studies differ from ours in several important aspects, including the experimental pipeline and configurations used, the set of LLMs evaluated, and the benchmark coverage. Prior work typically reports results for few models on one or two benchmarks (e.g., HumanEval or MBPP), whereas our study evaluates 20 SLMs across five distinct benchmarks. Due to these methodological differences, a direct quantitative comparison would not be reliable. Nevertheless, at a high level, comparing our results with existing results suggest that well-designed and carefully trained SLMs can achieve performance comparable to LLMs under specific benchmark conditions. This observation indicates that higher parameter count alone is not a dependable indicator of code generation quality, and that architectural choices, training procedures, and data curation play a critical role in determining practical performance.